# Analysis of Shuffling Beyond Pure Local Differential Privacy

### v1

@@ A.6. The proof of Theorem 5
Lemma A.10 (Band probability).
@@
(44) $\mathbb{E}_{Z_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\Bigr]\;\leq\;C\,\frac{c_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\,\varphi(t_{n}).$
@@
(45) $\mathbb{E}_{Z_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\Bigr]\;\leq\;C\,\frac{c_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\,\varphi(t_{n}),$
@@
(47) $\mathcal{D}^{\mathrm{blanket}}\;=\;\varphi\bigl(t_{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\varepsilon_{n}^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr),$
@@
(48) $=O\!\left(c_{n}\,\varepsilon_{n}^{2}n\right),$
@@
(49) $c_{n}\,\varepsilon_{n}^{2}n\;\lesssim\;\eta_{\mathrm{main}}.$
@@
(50) $c_{n}\;=\;\Theta\!\left(\frac{\eta_{\mathrm{main}}}{\varepsilon_{n}^{2}n}\right).$
@@
(51) $\Pr\bigl[|l_{\varepsilon_{n}}(Y)|>T\bigr]\;\leq\;C_{p}\,T^{-p},\qquad T>0.$
@@
(53) $=K^{\prime}_{\gamma,p}\,\varepsilon_{n}^{2}n^{5/2}\,\frac{\tau_{n}^{-p}}{\varphi(t_{n})}(1+o(1)).$
@@
(55) $\frac{E_{\mathrm{trunc}}(n;s_{n},w_{n}^{\mathrm{in}})}{\mathcal{D}^{\mathrm{blanket}}}\leq\frac{\eta_{\mathrm{trunc}}}{4},$
@@
(56) $\Pr\bigl[|E_{\mathrm{Mtrunc}}|>c_{n}/6\bigr]\;\leq\;\Pr[\perp]\;=\;E_{\mathrm{trunc}}(n;s_{n},w_{n}^{\mathrm{in}}).$
@@
(57) $\Pr\bigl[|E_{\mathrm{Mtrunc}}|>c_{n}/6\bigr]\;\leq\;\frac{\eta_{\mathrm{trunc}}}{4}\,\mathcal{D}^{\mathrm{blanket}}.$
@@
(58) $|\Delta_{\mathrm{ES}}|\leq K_{\mathrm{ES}}\,c_{n}\qquad\text{for all sufficiently large }n.$
@@
(59) $|\Delta_{\mathrm{ES}}|\leq\frac{c_{n}}{12}.$
@@
(60) $\Pr\left[R_{n}>c_{n}\right]\leq\exp\!\left(-c\,\frac{c_{n}^{2}}{nh_{n}^{2}}\right).$
@@
(61) $h_{n}:=\frac{c_{n}}{\sqrt{c\,n\,\Lambda_{h}}}.$
@@
(62) $\Pr\bigl[|E_{\mathrm{Mdisc}}|>c_{n}/6\bigr]\leq\Pr\bigl[|U_{n}|>c_{n}/12\bigr].$
@@
(63) $\Pr\bigl[|U_{n}|>c_{n}/12\bigr]\leq 2\exp\!\Bigl(-\kappa\,\frac{c_{n}^{2}}{nh_{n}^{2}}\Bigr),$
@@
(64) $\Pr\bigl[|E_{\mathrm{Mdisc}}|>c_{n}/6\bigr]\leq\frac{\eta_{\mathrm{disc}}}{4}\,\mathcal{D}^{\mathrm{blanket}}.$
@@
(65) $\Pr\bigl[|S_{n-1}^{\mathrm{di}}-\mu_{n}^{\mathrm{di}}|>u\bigr]\leq 2\exp\!\left(-\frac{u^{2}}{2\bigl(v_{n}+\tau_{n}u/3\bigr)}\right).$
@@
(67) $\Pr\bigl[|S_{n-1}^{\mathrm{di}}-\mu_{n}^{\mathrm{di}}|>u_{n}\bigr]\leq 2\exp(-c_{2}\Lambda_{w}).$
@@
(68) $w_{n}^{\mathrm{out}}:=2u_{n}+h_{n}=\Theta\bigl(\sqrt{v_{n}\Lambda_{w}}\bigr)=\Theta\bigl(\sqrt{n\Lambda_{w}}\bigr).$
@@
(69) $E_{\mathrm{alias}}(n;h_{n},s_{n},w_{n}^{\mathrm{in}},w_{n}^{\mathrm{out}}):=\Pr\bigl[\text{aliasing occurs}\bigr]\leq\Pr\bigl[|S_{n-1}^{\mathrm{di}}-\mu_{n}^{\mathrm{di}}|>u_{n}\bigr].$
@@
(70) $\delta_{\mathrm{alias}}\;\leq\;C_{\mathrm{alias},1}\,\eta_{\mathrm{alias}}\,\mathcal{D}^{\mathrm{blanket}}$
@@
(71) $w_{n}^{\mathrm{out}}=\Theta(\sqrt{n\log n}).$
@@
(72) $\Pr\bigl[|E_{\mathrm{Malias}}|>c_{n}/6\bigr]\leq\Pr\bigl[|E_{\mathrm{Malias}}|>0\bigr]\leq E_{\mathrm{alias}}(n;h_{n},s_{n},w_{n}^{\mathrm{in}},w_{n}^{\mathrm{out}}).$
@@
(73) $\Pr\bigl[|E_{\mathrm{Malias}}|>c_{n}/6\bigr]\leq\frac{\eta_{\mathrm{alias}}}{4}\,\mathcal{D}^{\mathrm{blanket}}.$
@@
(74) $\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]=\Pr\bigl[m_{n}(Z_{1})-\eta_{n}\left(\frac{3c_{n}}{2}\right)<T_{n-1}\leq m_{n}(Z_{1})+\eta_{n}\left(\frac{3c_{n}}{2}\right)\bigr].$
@@
(75) $|m|\leq C_{1}\sqrt{\log n},\qquad|\eta|\leq C_{1}/\sqrt{n},$
@@
(76) $\Pr\left[m-\eta<T_{n-1}\leq m+\eta\right]\;\leq\;C_{2}\,\eta\,\varphi(m).$
@@
(77) $|m|\leq C_{1}\sqrt{\log n},\qquad|\eta|\leq\frac{C_{1}}{\sqrt{n}}$
@@
(80) $\sum_{j=1}^{s-2}|I_{j,n}|\leq C_{2}^{\prime}\,\eta\,\varphi(m)\,n^{-1/2}(\log n)^{K}\leq C_{3}^{\prime}\,\eta\,\varphi(m),$
@@
(82) $\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\mathbf{1}\{A_{n}\}\leq C_{2}C_{5}\,\eta_{n}\left(\frac{3c_{n}}{2}\right)\,\varphi(t_{n})\mathbf{1}\{A_{n}\}.$
@@ Appendix B The Computation of Shuffle Indices
B.1. Generalized Gaussian Mechanisms
@@
Figure 2. Shuffle indices for generalized Gaussian mechanisms with different shape parameters $\beta$.
@@
For the upper shuffle index, we instead take the reference distribution to be one of the local randomizers $\mathcal{R}_{x}$ as in Su et al. (Su et al., 2025), and only the variance is needed in our analysis. From the definition, the variance is
(83) $\sigma_{x,\beta}^{2}(x_{1},x_{1}^{\prime}):=\int_{-\infty}^{\infty}\frac{\bigl(\phi_{x_{1}}^{\beta}(y)-\phi_{x_{1}^{\prime}}^{\beta}(y)\bigr)^{2}}{\phi_{x}^{\beta}(y)}\,dy\;=\;\alpha\frac{\beta}{2\alpha\,\Gamma(1/\beta)}\,F_{\beta}(a_{1},a_{1}^{\prime}),$

### v2

@@ A.6 The proof of Theorem 5
Remark A.9 (Band probability).
@@
$\mathbb{E}_{Z_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\Bigr]\;\leq\;C\,\frac{c_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\,\varphi(t_{n}).$ (44)
@@
$\mathbb{E}_{Z_{1}\sim\mathcal{R}_{x}}\Bigl[\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\Bigr]\;\leq\;C\,\frac{c_{n}}{\sigma_{\varepsilon_{n}}\sqrt{n-1}}\,\varphi(t_{n}),$ (45)
@@
$\mathcal{D}^{\mathrm{blanket}}\;=\;\varphi\bigl(t_{n}\bigr)\,\left(\frac{1}{\chi^{3}}\,\frac{1}{\varepsilon_{n}^{2}n^{3/2}}\right)\bigl(1+o(1)\bigr),$ (47)
@@
$=O\!\left(c_{n}\,\varepsilon_{n}^{2}n\right),$ (48)
@@
$c_{n}\,\varepsilon_{n}^{2}n\;\lesssim\;\eta_{\mathrm{main}}.$ (49)
@@
$c_{n}\;=\;\Theta\!\left(\frac{\eta_{\mathrm{main}}}{\varepsilon_{n}^{2}n}\right).$ (50)
@@
$\Pr\bigl[|l_{\varepsilon_{n}}(Y)|>T\bigr]\;\leq\;C_{p}\,T^{-p},\qquad T>0.$ (51)
@@
$=K^{\prime}_{\gamma,p}\,\varepsilon_{n}^{2}n^{5/2}\,\frac{\tau_{n}^{-p}}{\varphi(t_{n})}(1+o(1)).$ (53)
@@
$\frac{E_{\mathrm{trunc}}(n;s_{n},w_{n}^{\mathrm{in}})}{\mathcal{D}^{\mathrm{blanket}}}\leq\frac{\eta_{\mathrm{trunc}}}{4},$ (55)
@@
$\Pr\bigl[|E_{\mathrm{Mtrunc}}|>c_{n}/6\bigr]\;\leq\;\Pr[\perp]\;=\;E_{\mathrm{trunc}}(n;s_{n},w_{n}^{\mathrm{in}}).$ (56)
@@
$\Pr\bigl[|E_{\mathrm{Mtrunc}}|>c_{n}/6\bigr]\;\leq\;\frac{\eta_{\mathrm{trunc}}}{4}\,\mathcal{D}^{\mathrm{blanket}}.$ (57)
@@
$|\Delta_{\mathrm{ES}}|\leq K_{\mathrm{ES}}\,c_{n}\qquad\text{for all sufficiently large }n.$ (58)
@@
$|\Delta_{\mathrm{ES}}|\leq\frac{c_{n}}{12}.$ (59)
@@
$\Pr\left[R_{n}>c_{n}\right]\leq\exp\!\left(-c\,\frac{c_{n}^{2}}{nh_{n}^{2}}\right).$ (60)
@@
$h_{n}:=\frac{c_{n}}{\sqrt{c\,n\,\Lambda_{h}}}.$ (61)
@@
$\Pr\bigl[|E_{\mathrm{Mdisc}}|>c_{n}/6\bigr]\leq\Pr\bigl[|U_{n}|>c_{n}/12\bigr].$ (62)
@@
$\Pr\bigl[|U_{n}|>c_{n}/12\bigr]\leq 2\exp\!\Bigl(-\kappa\,\frac{c_{n}^{2}}{nh_{n}^{2}}\Bigr),$ (63)
@@
$\Pr\bigl[|E_{\mathrm{Mdisc}}|>c_{n}/6\bigr]\leq\frac{\eta_{\mathrm{disc}}}{4}\,\mathcal{D}^{\mathrm{blanket}}.$ (64)
@@
$\Pr\bigl[|S_{n-1}^{\mathrm{di}}-\mu_{n}^{\mathrm{di}}|>u\bigr]\leq 2\exp\!\left(-\frac{u^{2}}{2\bigl(v_{n}+\tau_{n}u/3\bigr)}\right).$ (65)
@@
$\Pr\bigl[|S_{n-1}^{\mathrm{di}}-\mu_{n}^{\mathrm{di}}|>u_{n}\bigr]\leq 2\exp(-c_{2}\Lambda_{w}).$ (67)
@@
$w_{n}^{\mathrm{out}}:=2u_{n}+h_{n}=\Theta\bigl(\sqrt{v_{n}\Lambda_{w}}\bigr)=\Theta\bigl(\sqrt{n\Lambda_{w}}\bigr).$ (68)
@@
$E_{\mathrm{alias}}(n;h_{n},s_{n},w_{n}^{\mathrm{in}},w_{n}^{\mathrm{out}}):=\Pr\bigl[\text{aliasing occurs}\bigr]\leq\Pr\bigl[|S_{n-1}^{\mathrm{di}}-\mu_{n}^{\mathrm{di}}|>u_{n}\bigr].$ (69)
@@
$\delta_{\mathrm{alias}}\;\leq\;C_{\mathrm{alias},1}\,\eta_{\mathrm{alias}}\,\mathcal{D}^{\mathrm{blanket}}$ (70)
@@
$w_{n}^{\mathrm{out}}=\Theta(\sqrt{n\log n}).$ (71)
@@
$\Pr\bigl[|E_{\mathrm{Malias}}|>c_{n}/6\bigr]\leq\Pr\bigl[|E_{\mathrm{Malias}}|>0\bigr]\leq E_{\mathrm{alias}}(n;h_{n},s_{n},w_{n}^{\mathrm{in}},w_{n}^{\mathrm{out}}).$ (72)
@@
$\Pr\bigl[|E_{\mathrm{Malias}}|>c_{n}/6\bigr]\leq\frac{\eta_{\mathrm{alias}}}{4}\,\mathcal{D}^{\mathrm{blanket}}.$ (73)
@@
$\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]=\Pr\bigl[m_{n}(Z_{1})-\eta_{n}\left(\frac{3c_{n}}{2}\right)<T_{n-1}\leq m_{n}(Z_{1})+\eta_{n}\left(\frac{3c_{n}}{2}\right)\bigr].$ (74)
@@
$|m|\leq C_{1}\sqrt{\log n},\qquad|\eta|\leq C_{1}/\sqrt{n},$ (75)
@@
$\Pr\left[m-\eta<T_{n-1}\leq m+\eta\right]\;\leq\;C_{2}\,\eta\,\varphi(m).$ (76)
@@
$|m|\leq C_{1}\sqrt{\log n},\qquad|\eta|\leq\frac{C_{1}}{\sqrt{n}}$ (77)
@@
$\sum_{j=1}^{s-2}|I_{j,n}|\leq C_{2}^{\prime}\,\eta\,\varphi(m)\,n^{-1/2}(\log n)^{K}\leq C_{3}^{\prime}\,\eta\,\varphi(m),$ (80)
@@
$\Pr\bigl[|S_{n-1}+Z_{1}+\Delta_{\mathrm{ES}}|<\tfrac{3}{2}c_{n}\,\big|\,Z_{1}\bigr]\mathbf{1}\{A_{n}\}\leq C_{2}C_{5}\,\eta_{n}\left(\frac{3c_{n}}{2}\right)\,\varphi(t_{n})\mathbf{1}\{A_{n}\}.$ (82)
@@ Appendix B The Computation of Shuffle Indices
B.1 Generalized Gaussian Mechanisms
@@
Figure 2: Shuffle indices for generalized Gaussian mechanisms with different shape parameters $\beta$.
@@
For the upper shuffle index, we instead take the reference distribution to be one of the local randomizers $\mathcal{R}_{x}$ as in Su et al. Su et al. (2025), and only the variance is needed in our analysis. From the definition, the variance is
$\sigma_{x,\beta}^{2}(x_{1},x_{1}^{\prime}):=\int_{-\infty}^{\infty}\frac{\bigl(\phi_{x_{1}}^{\beta}(y)-\phi_{x_{1}^{\prime}}^{\beta}(y)\bigr)^{2}}{\phi_{x}^{\beta}(y)}\,dy\;=\;\alpha\frac{\beta}{2\alpha\,\Gamma(1/\beta)}\,F_{\beta}(a_{1},a_{1}^{\prime}),$ (83)